# PoGDiff: Product-of-Gaussians Diffusion Models for Imbalanced Text-to-Image Generation

### v2

@@ 1 Introduction
This limitation is true even for finetuning large diffusion models pretrained on large-scale datasets like LAION-5B [31], e.g., Stable Diffusion [28]. Imagine an imbalanced dataset consisting of employees in a small company, senior employees might have more photos available, while new employees only have a very limited number of them. Since none of the employees appear in the LAION-5B dataset, generating photos of them require finetuning the Stable Diffusion model. Unfortunately, finetuning the model on such an imbalanced dataset might enable the model to generate accurate images for the majority group (i.e., senior employees), but it will perform poorly for the minority group (i.e., new employees).
@@
  • We identify the problem of imbalanced text-to-image generation (IT2I) and introduce the first general diffusion model, dubbed Product-of-Gaussians Diffusion Models (PoGDiff), for addressing this problem.
@@ 2 Related Work
Long-Tailed Recognition. Addressing the challenges posed by long-tailed data distributions has been a critical area of research in machine learning, for both classification and regression problems. Traditional methods, such as re-sampling and re-weighting techniques, have been used to mitigate class imbalances by either over-sampling minority classes or assigning higher weights to them during training [7, 9, 37, 3, 4]. Such algorithms fail to measure the distance in continuous label space and fall short in handling high-dimensional data (e.g., images and text). Deep imbalanced regression methods [39, 27, 8, 15, 38] address this challenge by reweighting the data using the effective label density during representation learning. However, all methods above are designed for recognition tasks such as classification and regression, and are therefore not applicable to our generation task.
Diffusion Models Related to Long-Tailed Data. There are also works that related to both diffusion models and long-tailed data. They aim at improving generation robustness using noisy label [22], improving fairness in image generation [32], and improving classification accuracy using diffusion models [40]. However, these works have different goals and therefore are not applicable to our setting.
Most relevant to our work is Class Balancing Diffusion Model (CBDM) [26], which uses a distribution adjustment regularizer that enhances tail-class generation based on the model’s predictions for the head class. It improves the quality of long-tailed generation by assuming one-hot conditional labels (i.e., classification-based settings). However, this assumption does not generalize to the modern setting where image generation is usually conditioned on free-form text prompts. As a result, when adapted to the free-form setting, they often fail to model the similarity among different text prompts, leading to suboptimal generation performance in minority data (as verified by empirical results in Sec. 4).
@@ 3.1 Preliminaries
$q\left({\bf x}_{1:T}|{\bf x}_{0}\right)$ $=\prod\nolimits_{t=1}^{T}q\left({\bf x}_{t}|{\bf x}_{t-1}\right)$
$q\left({\bf x}_{t}|{\bf x}_{0}\right)$ $=\mathcal{N}\left({\bf x}_{t};\sqrt{\bar{\alpha}_{t}}{\bf x}_{0},(1-\bar{\alpha}_{t}){\bf I}\right)$
@@
This can be simplified to:
$L_{DM}$ $=\mathbb{E}_{{\bf x}_{0}={\bf x},\epsilon\sim\mathcal{N}(0,{\bf I}),t}\left[\|\epsilon-\epsilon_{\theta}({\bf x}_{t},t,{\bf y})\|_{2}^{2}\right].$
Latent diffusion models (LDMs) [28] are diffusion models that perform the entire diffusion and denoising process in a lower-dimensional latent space. LDMs first learn an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, which are then frozen during subsequent training of the diffusion models. The corresponding objective is then simplified to:
$L_{LDM}=\mathbb{E}_{{\bf z}_{0}=\mathcal{E}({\bf x}),\epsilon\sim\mathcal{N}(0,{\bf I}),t}\left[\|\epsilon-\epsilon_{\theta}({\bf z}_{t},t,{\bf y})\|_{2}^{2}\right]$
In this paper, we use Stable Diffusion (LDM) [28] as our backbone model. Since our method works for both the vanilla DMs and LDMs, for clarity, we use the notation ${\bf x}$ instead of ${\bf z}$, as the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are fixed during fine-tuning.
@@ 3.2.1 Main Idea
Typical diffusion models minimize the KL divergence between the predicted distribution $p_{\theta}({\bf x}_{t-1}|{\bf x}_{t},{\bf y})=\mathcal{N}(\epsilon_{\theta}({\bf x}_{t},t,{\bf y}),\lambda_{{\bf y}}^{-1}{\bf I})$ and the ground-truth distribution $q({\bf x}_{t-1}|{\bf x}_{t},{\bf x}_{0},{\bf y})=\mathcal{N}(\epsilon,\lambda_{t}^{-1}{\bf I})$ at each time step $t$ during the backward denoising process. Here, $\lambda_{{\bf y}}$ and $\lambda_{t}$ represent the precision. In contrast, our PoGDiff replaces the ground-truth target with a Product of Gaussians (PoG), and instead minimize the following KL divergence (for each $t$)
$\mathcal{L}_{t-1}^{\textrm{PoGDiff}}=D_{KL}\big{(}$ $q\left({\bf x}_{t-1}|{\bf x}_{t},{\bf x}_{0},{\bf y}\right)\circ p_{\theta}\left({\bf x}_{t-1}|{\bf x}_{t},{\bf y}^{\prime}\right)$
$\big{\|}p_{\theta}\left({\bf x}_{t-1}|{\bf x}_{t},{\bf y}\right)\big{)},$ (1)
@@
$\mathcal{N}(\mu_{1},\lambda_{1}^{-1})\circ\mathcal{N}(\mu_{2},\lambda_{2}^{-1})$
$\quad=\mathcal{N}\left(\dfrac{\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2}}{\lambda_{1}+\lambda_{2}},(\lambda_{1}+\lambda_{2})^{-1}\right)$
$\quad\triangleq\mathcal{N}\left(\mu_{\textrm{PoG}},\lambda_{\textrm{PoG}}^{-1}\right),$ (2)
@@ 3.2.2 Theoretical Analysis and Algorithmic Design
$\mathcal{L}_{t-1}^{\textrm{PoGDiff}}$ $\leq\mathbb{E}_{q}\Big{[}\mathcal{A}(\lambda_{t})\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}$
$\quad+\mathcal{A}(\lambda_{{\bf y}^{\prime}})\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]}+C$ (4)
@@
The proof is available in the Appendix A. Eqn. (4) in Proposition 3.1 provides a upper bound for the KL divergence (Eqn. (1)) we aim to minimize.
@@
$\mathcal{L}_{\textrm{simple}}^{\textrm{PoGDiff}}$ $=\mathbb{E}_{{\bf x}_{0}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\Big{[}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}$
$\quad+\dfrac{\lambda_{{\bf y}^{\prime}}}{\lambda_{t}}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]}.$ (5)
@@
$\mathcal{L}_{\textrm{simple}}^{\textrm{PoGDiff}}$ $=\mathbb{E}_{{\bf x}_{0}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\Big{[}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}$
$\quad+\psi\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]}.$ (6)
@@ 3.2.3 Computing the Similarity $\psi$
Next, we discuss the choice of $\psi$ in Eqn. (6). Given a image-text dataset $\mathcal{D}$, the similarities between each image-text pair need to be considered in two parts:
@@
3.2.4 Final Objective Function
Figure 4: Overview of label distribution for four IT2I datasets. The x-axis corresponds to the identities (i.e., people or individuals).
where $s\in\{s_{j}\}_{j=1}^{k}$ denotes the cosine similarity sampled according to the weights $\{w_{j}\}_{j=1}^{k}$ in Eqn. (8), $\mathbbm{1}\left[\cdot\right]$ denotes the indicator function, and $\mathcal{I}(\cdot)$ retrieves the class/identity of the current input image; $\mathbbm{1}\left[\mathcal{I}({\bf x})\neq\mathcal{I}({\bf x}^{\prime})\right]=0$ if ${\bf x}$ and ${\bf x}^{\prime}$ are two photos of the same person (e.g., Albert Einstein), and $\mathbbm{1}\left[\mathcal{I}({\bf x})\neq\mathcal{I}({\bf x}^{\prime})\right]=1$ if ${\bf x}$ and ${\bf x}^{\prime}$ are photos of two different persons (e.g., ${\bf x}$ is Einstein and ${\bf x}^{\prime}$ Reagan). $a_{1},a_{2}$ are hyperparameters that control the scale of the similarities. For example, if the cosine similarity ($s$) between $x$ and $x^{\prime}$ is 0.4, and $a_{1}=a_{2}=1$: if $x$ and $x^{\prime}$ are of the same person, the image similarity will be $0.4^{1}$, whereas if $x$ and $x^{\prime}$ are not of the same person, the image similarity will be $0.4^{2}$, which is smaller. The intuition is to compute the image similarity according to both the image content similarity, i.e., $s$, and identity similarity, i.e., $\mathcal{I}({\bf x})$ and $\mathcal{I}({\bf x}^{\prime})$.
@@ 3.2.4 Final Objective Function
$\mathcal{L}_{\textrm{final}}^{\textrm{PoGDiff}}$ $=\mathbb{E}_{{\bf x}_{0}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\Big{[}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}$
$\quad+\psi\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]},$ (13)
@@ 4.1 Experimental Setup
Datasets. To demonstrate the effectiveness of PoGDiff in terms of both accuracy and quality, we evaluate our method on two widely used imbalanced datasets, i.e., AgeDB-IT2I [21] and DigiFace-IT2I [2].
Note that our method is designed for fine-tuning. Therefore our setup does not require large-scale, long-tailed human datasets. Instead, we sample from these datasets, as long as they meet the following criteria: (1) the dataset must be long-tailed, (2) traditional methods must fail to recognize the minority classes, and (3) there must be a distinguishable difference between the majority and minority classes (e.g., we prefer visual distinctions between the two groups to better highlight the impact of our method). Fig. 4 shows the label density distribution of these datasets, and their level of imbalance (see Appendix G.4 for details on data sparsity).
Table 1: Performance based on FID score.
Datasets AgeDB-IT2I DigiFace-IT2I VGGFace-IT2I
Size Small Medium Large Large Large
Metric FID $\downarrow$
Shot All Few All Few All Few All Few All Few
Vanilla 14.88 13.72 12.87 12.56 7.67 11.67 7.18 12.23 7.59 12.08
CBDM 14.72 14.13 11.63 11.59 7.18 11.12 6.96 12.72 7.23 11.91
T2H 14.85 13.66 12.79 12.52 7.61 11.64 7.14 12.22 7.34 12.02
PoGDiff (Ours) 14.15 12.88 10.89 10.64 6.03 10.16 6.84 11.21 6.29 10.97
Table 2: Performance based on DINO score.
Datasets AgeDB-IT2I DigiFace-IT2I VGGFace-IT2I
Size Small Medium Large Large Large
Metric DINO (cosine similarity) scores $\uparrow$
Shot All Few All Few All Few All Few All Few
Vanilla 0.42 0.37 0.39 0.28 0.34 0.25 0.42 0.36 0.41 0.29
CBDM 0.54 0.09 0.38 0.11 0.41 0.26 0.34 0.16 0.46 0.22
T2H 0.43 0.39 0.42 0.29 0.37 0.26 0.44 0.36 0.42 0.28
PoGDiff (Ours) 0.77 0.73 0.69 0.56 0.66 0.52 0.64 0.49 0.69 0.55
Table 3: Performance on AgeDB-IT2I based on human evaluation. The evaluation is a binary decision: Image is either judged as representing the same individual (score 1.0) or not (score 0.0).
Datasets AgeDB-IT2I VGGFace-IT2I
Size Small Medium Large Large
Metric Human Score $\uparrow$
Shot All Few All Few All Few All Few
Vanilla 0.50 0.00 0.66 0.32 0.60 0.20 0.62 0.16
CBDM 0.50 0.00 0.44 0.08 0.56 0.12 0.54 0.10
T2H 0.50 0.00 0.66 0.32 0.60 0.20 0.62 0.16
PoGDiff (Ours) 1.00 1.00 0.96 0.92 0.84 0.68 0.78 0.64
Table 4: Performance on AgeDB-IT2I based on GPT-4o evaluation. The scores are from 0 to 10, with higher scores indicating the individual resembles the well-known person.
Datasets AgeDB-IT2I VGGFace-IT2I
Size Small Medium Large Large
Metric GPT-4o Evaluation $\uparrow$
Shot All Few All Few All Few All Few
Vanilla 5.20 3.20 4.30 2.90 4.90 3.60 4.50 2.90
CBDM 4.50 1.10 1.30 1.00 3.10 1.70 2.80 1.30
T2H 5.50 3.10 4.60 3.00 4.70 3.90 4.60 3.10
PoGDiff (Ours) 9.10 8.40 8.80 8.20 8.50 8.00 8.20 7.60
@@
DigiFace-IT2I-L: DigiFace-IT2I-L is derived from the DigiFace dataset [2]. It contains $985$ images across $179$ identities, where each majority class consists of $30$ images and each minority class consists of $2$ images. We use a process similar to AgeDB-IT2I to collect text-image pairs, forming this DigiFace-IT2I dataset.
Figure 5: Example generated images from different methods. Our PoGDiff outperforms the baselines in both generation accuracy and quality. Regarding the ground truth (GT), the training set for the minority class (left two columns) contains only $2$ images per individual while the majority class has more than $10$ samples per individual.
@@
VGGFace-IT2I-L: VGGFace-IT2I-L is a subset from VGGFace2 [5]. It contains $1933$ images across $193$ identities, where each majority class consists of $49$ images and each minority class consists of $2$ images.
Baselines. We employ Stable Diffusion v1.5 [28] as the backbone diffusion model, and incorporate two recent works for handling classification-based imbalanced diffusion model into comparison. Details for each baselines are in the Appendix B.
@@
Evaluation Protocols and Metrics. We use three types of evaluation metrics: generation quality, generation accuracy, and generation diversity.
For general text-to-image generation quality, we report the widely used Fréchet Inception Distance (FID) score [10]. Unlike the traditional FID score, which uses Inception-v3 [36] as the feature extractor, we used a pre-trained face recognition model instead; since our goal is to evaluate the ability to recognize humans, we need to capture facial features rather than general features. For each identity, we collect all images from the original AgeDB or DigiFace datasets as the true image set. Then, In all-shot evaluation, for AgeDB-IT2I-S and AgeDB-IT2I-M, we generate $100$ images per identity as the fake image set, and for AgeDB-IT2I-L and DigiFace-IT2I-L, we generate $20$ images per identity as the fake image set. In few-shot evaluation, we generate $500$ images per identity as the fake image set. For all generations, we employ the DDIM sampling technique [34] with $50$ steps. The prompt used during generation is “An image of {p}.” where “p” is the name of the identity (e.g., Albert Einstein).
To assess generation accuracy, we use 10 different seeds to sample 10 images for each minority class. We then gather feedback from both the GPT-4o model [1] and human evaluators to score the accuracy of identity recognition. Additionally, we employ a pre-trained DINO model [6] for calculating DINO score for image similarities. More details about the evaluation process including prompts we used are in the Appendix C.
@@
  • Analysis. This metric evaluates the generational diversity of a model. For example, if the training dataset contains two distinct images of Einstein, $x_{1}$ and $x_{2}$, and a model generates only images resembling $x_{1}$, the gRecall in this case would be $0.5$. While the model may achieve high accuracy in terms of facial identity ( Table 3 and  Table 4), it falls short in diversity because it fails to generate images resembling $x_{2}$. In contrast, if a model generates images that cover both $x_{1}$ and $x_{2}$ the gRecall for this ID will be $1$; for instance, if the model generates 10 images for Einstein, where 6 of them resemble $x_{1}$ and 4 of them resemble $x_{2}$, the gRecall would be $1$, indicating high diversity and coverage.

### v3

@@ 1 Introduction
This limitation is true even for finetuning large diffusion models pretrained on large-scale datasets like LAION-5B [16], e.g., Stable Diffusion [7]. Imagine an imbalanced dataset consisting of employees in a small company, senior employees might have more photos available, while new employees only have a very limited number of them. Since none of the employees appear in the LAION-5B dataset, generating photos of them requires finetuning the Stable Diffusion model. Unfortunately, finetuning the model on such an imbalanced dataset might enable the model to generate accurate images for the majority group (i.e., senior employees), but it will perform poorly for the minority group (i.e., new employees).
@@
  • We identify the problem of imbalanced text-to-image generation (IT2I) and introduce the first general diffusion model, dubbed Product-of-Gaussians Diffusion Models (PoGDiff), to address this problem.
@@ 2 Related Work
Long-Tailed Recognition. Addressing the challenges posed by long-tailed data distributions has been a critical area of research in machine learning, for both classification and regression problems. Traditional methods, such as re-sampling and re-weighting techniques, have been used to mitigate class imbalances by either over-sampling minority classes or assigning higher weights to them during training [17, 18, 19, 20, 21]. Such algorithms fail to measure the distance in continuous label space and fail to handle high-dimensional data (e.g., images, and text). Deep imbalanced regression methods [22, 23, 24, 25, 26] address this challenge by reweighting the data using the effective label density during representation learning. However, all methods above are designed for recognition tasks such as classification and regression and are therefore not applicable to our generation task.
Diffusion Models Related to Long-Tailed Data. There are also works related to both diffusion models and long-tailed data. They aim at improving generation robustness using feature engineering [27, 28, 29], feature augmentation [30], noisy label [31], improving fairness in image generation [32, 33, 34], and improving classification accuracy using diffusion models [30]. However, these works have different goals and, therefore, are not applicable to our setting.
Most relevant to our work is the Class Balancing Diffusion Model (CBDM) [15], which uses a distribution adjustment regularizer that enhances tail-class generation based on the model’s predictions for the head class. It improves the quality of long-tailed generation by assuming one-hot conditional labels (i.e., classification-based settings). However, this assumption does not generalize to the modern setting where image generation is usually conditioned on free-form text prompts. As a result, when adapted to the free-form setting, they often fail to model the similarity among different text prompts, leading to suboptimal generation performance in minority data (as verified by empirical results in Sec. 4).
@@ 3.1 Preliminaries
$q\left({\bf x}_{1:T}|{\bf x}_{0}\right)=\prod\nolimits_{t=1}^{T}q\left({\bf x}_{t}|{\bf x}_{t-1}\right),\quad q\left({\bf x}_{t}|{\bf x}_{0}\right)=\mathcal{N}\left({\bf x}_{t};\sqrt{\bar{\alpha}_{t}}{\bf x}_{0},(1-\bar{\alpha}_{t}){\bf I}\right)$
@@
This can be simplified to (i.e., Eqn. (1) in Rombach et al. [7]):
$L_{DM}$ $=\mathbb{E}_{{\bf x},\epsilon\sim\mathcal{N}(0,{\bf I}),t}\left[\|\epsilon-\epsilon_{\theta}({\bf x}_{t},t,{\bf y})\|_{2}^{2}\right].$
Latent diffusion models (LDMs) [7] are diffusion models that perform the entire diffusion and denoising process in a lower-dimensional latent space. LDMs first learn an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, which are then frozen during subsequent training of the diffusion models. The corresponding objective is then simplified to (i.e., Eqn. (2) in Rombach et al. [7]):
$L_{LDM}=\mathbb{E}_{\mathcal{E}({\bf x}),\epsilon\sim\mathcal{N}(0,{\bf I}),t}\left[\|\epsilon-\epsilon_{\theta}({\bf z}_{t},t,{\bf y})\|_{2}^{2}\right]$
In this paper, we use Stable Diffusion (SD) [7] as our backbone model. Since our method works for both the vanilla DMs and LDMs, for clarity, we use the notation ${\bf x}$ instead of ${\bf z}$, as the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ are fixed during fine-tuning.
@@ 3.2.1 Main Idea
Typical diffusion models minimize the KL divergence between the predicted distribution $p_{\theta}({\bf x}_{t-1}|{\bf x}_{t},{\bf y})=\mathcal{N}({{\bf x}_{\theta}}({\bf x}_{t},t,{\bf y}),\lambda_{{\bf y}}^{-1}{\bf I})$ and the ground-truth distribution $q({\bf x}_{t-1}|{\bf x}_{t},{\bf x}_{0},{\bf y})=\mathcal{N}({{\bf x}_{t-1}},\lambda_{t}^{-1}{\bf I})$ at each time step $t$ during the backward denoising process. Here, $\lambda_{{\bf y}}$ and $\lambda_{t}$ represent the precision. In contrast, our PoGDiff replaces the ground-truth target with a Product of Gaussians (PoG) and instead minimizes the following KL divergence (for each $t$)
$\mathcal{L}_{t-1}^{\textrm{PoGDiff}}=D_{KL}\big{(}q\left({\bf x}_{t-1}|{\bf x}_{t},{\bf x}_{0},{\bf y}\right)\circ p_{\theta}\left({\bf x}_{t-1}|{\bf x}_{t},{\bf y}^{\prime}\right)\big{\|}p_{\theta}\left({\bf x}_{t-1}|{\bf x}_{t},{\bf y}\right)\big{)},$ (1)
@@
$\mathcal{N}(\mu_{1},\lambda_{1}^{-1})\circ\mathcal{N}(\mu_{2},\lambda_{2}^{-1})=\mathcal{N}\left(\dfrac{\lambda_{1}\mu_{1}+\lambda_{2}\mu_{2}}{\lambda_{1}+\lambda_{2}},(\lambda_{1}+\lambda_{2})^{-1}\right)\triangleq\mathcal{N}\left(\mu_{\textrm{PoG}},\lambda_{\textrm{PoG}}^{-1}\right),$ (2)
@@ 3.2.2 Theoretical Analysis and Algorithmic Design
$\mathcal{L}_{t-1}^{\textrm{PoGDiff}}\leq\mathbb{E}_{q}\Big{[}\mathcal{A}(\lambda_{t})\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}+\mathcal{A}(\lambda_{{\bf y}^{\prime}})\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]}+C$ (4)
@@
The proof is available in  Appendix A. Eqn. (4) in Proposition 3.1 provides a upper bound for the KL divergence (Eqn. (1)) we aim to minimize.
@@
$\mathcal{L}_{\textrm{simple}}^{\textrm{PoGDiff}}=\mathbb{E}_{{\bf x}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\left[\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}+\dfrac{\lambda_{{\bf y}^{\prime}}}{\lambda_{t}}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\right].$ (5)
@@
$\mathcal{L}_{\textrm{simple}}^{\textrm{PoGDiff}}=\mathbb{E}_{{\bf x}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\Big{[}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}+\psi\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]}.$ (6)
@@ 3.2.3 Computing the Similarity $\psi$
Next, we discuss the choice of $\psi$ in Eqn. (6). Given an image-text dataset $\mathcal{D}$, the similarities between each image-text pair need to be considered in two parts:
@@
where $s\in\{s_{j}\}_{j=1}^{k}$ denotes the cosine similarity sampled according to the weights $\{w_{j}\}_{j=1}^{k}$ in Eqn. (8), $\mathbbm{1}\left[\cdot\right]$ denotes the indicator function, and $\mathcal{I}(\cdot)$ retrieves the class/identity of the current input image; $\mathbbm{1}\left[\mathcal{I}({\bf x})\neq\mathcal{I}({\bf x}^{\prime})\right]=0$ if ${\bf x}$ and ${\bf x}^{\prime}$ are two photos of the same person (e.g., Albert Einstein), and $\mathbbm{1}\left[\mathcal{I}({\bf x})\neq\mathcal{I}({\bf x}^{\prime})\right]=1$ if ${\bf x}$ and ${\bf x}^{\prime}$ are photos of two different persons (e.g., ${\bf x}$ is Einstein and ${\bf x}^{\prime}$ Reagan). More details on $\mathcal{I}(\cdot)$ can be found in Appendix J.7. $a_{1},a_{2}$ are hyperparameters that control the scale of the similarities. For example, if the cosine similarity ($s$) between $x$ and $x^{\prime}$ is 0.4, and $a_{1}=a_{2}=1$: if $x$ and $x^{\prime}$ are of the same person, the image similarity will be $0.4^{1}$, whereas if $x$ and $x^{\prime}$ are not of the same person, the image similarity will be $0.4^{2}$, which is smaller. The intuition is to compute the image similarity according to both the image content similarity, i.e., $s$, and identity similarity, i.e., $\mathcal{I}({\bf x})$ and $\mathcal{I}({\bf x}^{\prime})$.
@@ 3.2.4 Final Objective Function
$\mathcal{L}_{\textrm{final}}^{\textrm{PoGDiff}}$ $=\mathbb{E}_{{\bf x}\sim q({\bf x}_{0}),\epsilon\sim\mathcal{N}({\bf 0},{\bf I}),t\sim\mathcal{U}(1,T)}\Big{[}\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon\right\|^{2}+\psi\left\|\epsilon_{\theta}({\bf x}_{t},{\bf y})-\epsilon_{\theta}({\bf x}_{t},{\bf y}^{\prime})\right\|^{2}\Big{]},$ (13)
@@ 4.1 Experimental Setup
Datasets. To demonstrate the effectiveness of PoGDiff in terms of both accuracy and quality, we evaluate our method on four widely used imbalanced datasets, i.e., AgeDB [35], DigiFace [36], VGGFace2 [37] and CIFAR-100-LT [38].
AgeDB-IT2I-L & AgeDB-IT2I-S: AgeDB-IT2I-L(arge) is constructed from the AgeDB dataset [35]. It consists of $976$ images across $223$ identities, with each majority class containing $30$ images and each minority class containing $2$ images. We also construct AgeDB-IT2I-S(mall), which contains $32$ images across $2$ identities, where each majority class consists of $30$ images and each minority class consists of $2$ images. Additionally, we construct AgeDB-IT2I-M(edium), and more details can be found in Appendix C.
@@
DigiFace-IT2I: DigiFace-IT2I is derived from the DigiFace dataset [36]. It contains $985$ images across $179$ identities, where each majority class consists of $30$ images and each minority class consists of $2$ images. We use a process similar to AgeDB-IT2I to collect text-image pairs, forming this DigiFace-IT2I dataset.
@@
VGGFace-IT2I: VGGFace-IT2I is a subset of VGGFace2 [37]. It contains $1933$ images across $193$ identities, where each majority class consists of $49$ images and each minority class consists of $2$ images.
CIFAR-100-IT2I: CIFAR-100-IT2I is constructed from CIFAR-100-LT dataset [38]. We set the imbalance ratio to 250, where the largest class contains $500$ images and the smallest class contains only $2$ images. The dataset consists of $9502$ images in total. During fine-tuning, to mitigate the dominance of majority classes, we downsample these classes in each epoch. Specifically, the most frequent class is reduced to $49$ images per epoch out of its original $500$ images.
Table 1: Performance based on FID score. Datasets (-IT2I) AgeDB DigiFace VGGFace CIFAR-100 Size Small Large Large Large Large Metric FID $\downarrow$ Shot All Few All Few All Few All Few All Few Vanilla 14.88 13.72 7.67 11.67 7.18 12.23 7.59 12.08 7.19 11.46 CBDM 14.72 14.13 7.18 11.12 6.96 12.72 7.23 11.91 7.26 11.94 T2H 14.85 13.66 7.61 11.64 7.14 12.22 7.34 12.02 7.10 11.39 PoGDiff (Ours) 14.15 12.88 6.03 10.16 6.84 11.21 6.29 10.97 6.24 9.41
Table 2: Performance based on DINO score. Datasets (-IT2I) AgeDB DigiFace VGGFace CIFAR-100 Size Small Large Large Large Large Metric DINO $\uparrow$ Shot All Few All Few All Few All Few All Few Vanilla 0.42 0.37 0.34 0.25 0.42 0.36 0.41 0.29 0.48 0.32 CBDM 0.54 0.09 0.41 0.26 0.34 0.16 0.46 0.22 0.52 0.28 T2H 0.43 0.39 0.37 0.26 0.44 0.36 0.42 0.28 0.45 0.30 PoGDiff (Ours) 0.77 0.73 0.66 0.52 0.64 0.49 0.69 0.55 0.73 0.61
Table 3: Performance based on human evaluation. The evaluation is a binary decision: the image is either judged as representing the same individual (score 1.0) or not (score 0.0). Datasets (-IT2I) AgeDB VGGFace CIFAR-100 Size Small Large Large Large Metric Human Score $\uparrow$ Shot All Few All Few All Few All Few Vanilla 0.50 0.00 0.60 0.20 0.62 0.16 0.72 0.30 CBDM 0.50 0.00 0.56 0.12 0.54 0.10 0.63 0.24 T2H 0.50 0.00 0.60 0.20 0.62 0.16 0.72 0.30 PoGDiff (Ours) 1.00 1.00 0.84 0.68 0.78 0.64 0.84 0.72
Table 4: Performance on AgeDB-IT2I based on GPT-4o evaluation. The scores are from 0 to 10, with higher scores indicating the individual resembles the well-known person. Datasets (-IT2I) AgeDB VGGFace CIFAR-100 Size Small Large Large Large Metric Human Score $\uparrow$ Shot All Few All Few All Few All Few Vanilla 5.20 3.20 4.90 3.60 4.50 2.90 6.00 3.20 CBDM 4.50 1.10 3.10 1.70 2.80 1.30 3.40 2.00 T2H 5.50 3.10 4.70 3.90 4.60 3.10 6.20 3.60 PoGDiff (Ours) 9.10 8.40 8.50 8.00 8.20 7.60 8.40 8.00
Baselines. We employ Stable Diffusion v1.5 [7] as the backbone diffusion model. As this is the first work to explore imbalanced text-to-image (IT2I) diffusion models with natural text prompts, we adapt the current state-of-the-art methods designed for long-tailed T2I diffusion models with one-hot text prompts to serve as baselines. The baselines are described below:
  • Vanilla: The Vanilla model simply fine-tunes a Stable Diffusion model without additional modifications.
  • CBDM: CBDM [15] is a Class Balancing Diffusion Model that incorporates a distribution adjustment regularizer during training. During fine-tuning, we sample an additional text embedding ${\bf y}^{\prime}$ from the entire fine-tuning dataset and apply the CBDM objective function. All hyperparameters are kept the same as in the original paper, with further details available in Qin et al. [15].
  • T2H: T2H [30] is a Long-Tailed Diffusion Model with Oriented Calibration. It is a feature augmentation method, but is not directly applicable to our setting. Specifically, T2H relies on the class frequency, which is not available in our experiments. We adapted this method to our settings by using the density for each text prompt embedding to serve as the class frequency in T2H [30].
@@
Evaluation Protocols and Metrics. We use three types of evaluation metrics: generation diversity, generation accuracy, and generation quality.
@@
  • Analysis. This metric evaluates the generational diversity of a model. For example, if the training dataset contains two distinct images of Einstein, $x_{1}$ and $x_{2}$, and a model generates only images resembling $x_{1}$, the gRecall, in this case, would be $0.5$. While the model may achieve high accuracy in terms of facial identity ( Table 4 and  Table 4), it falls short in diversity because it fails to generate images resembling $x_{2}$. In contrast, if a model generates images that cover both $x_{1}$ and $x_{2}$, the gRecall for this ID will be $1$; for instance, if the model generates 10 images for Einstein, where 6 of them resemble $x_{1}$, and 4 of them resemble $x_{2}$, the gRecall would be $1$, indicating high diversity and coverage.
To assess generation accuracy, we use 10 different seeds to sample 10 images for each minority class. We then gather feedback from both the GPT-4o model [39] and human evaluators to score the accuracy of identity recognition. Additionally, we employ a pre-trained DINO model [40] for calculating the DINO score for image similarities. More details about the evaluation process, including prompts we used, are in  Appendix E.
Figure 4: Example generated images from different methods. Our PoGDiff outperforms the baselines in both generation accuracy and quality. Regarding the ground truth (GT), the training set for the minority class (left two columns) contains only $2$ images per individual while the majority class has more than $8$ samples per individual.
For general text-to-image generation quality, we report the widely used Fréchet Inception Distance (FID) score [41]. For all the facial datasets, we use a pre-trained face recognition model as the feature extractor rather than traditional Inception-v3 [42]; since our goal is to evaluate the ability to recognize humans, we need to capture facial features rather than general features. For CIFAR-100-LT, we employ the original feature extractor (i.e., Inception-v3 [42]). More details are in Appendix E.